2. A student carried out a reaction between iron and hydrochloric acid to make some iron (II) chloride crystals. She started with 5.6g of iron and used an excess of hydrochloric acid.

a. Write a balanced chemical equation for the reaction!    [3]

b. How many moles of iron did he start off with?    [2]

c. What mass of iron (II) chloride should he have expected from this reaction?    [2]

d. He actually obtained 9.17g of the iron (II) chloride. What was his percentage yield? [2]


1
Expert's answer
2021-06-07T03:15:54-0400

a)Fe + 2HCl -->> FeCl2 + H2

b) Molar Mass of iron = 55.845

= 5.6/55.845

= 0.1 m

c) Molar Mass of FeCl2 = 126.751

= 126.751 × 0.1

= 12.71g

d) 12.71-9.17

= 3.54

= 3.54/12.71 × 100

= 27.85%